OpenStudy (natsent06):

Select the TRUE statement. A) (b + c)a = a + (b × c) B) (b + c)a = (a + b) × (a + c) C) Both A and B D) Neither A nor B

OpenStudy (igreen):

A is the only true statement.

OpenStudy (igreen):

If you input: a = 2 b = 3 c = 4 (b + c)a = 14 a + (b × c) = 14 (a + b) × (a + c) = 30